# UniCtrl: Improving the Spatiotemporal Consistency of Text-to-Video Diffusion Models via Training-Free Unified Attention Control

### v3

@@ Abstract
Keywords:
Video Diffusion Spatiotemporal Consistency Attention
@@ 1 Introduction
Diffusion Models (DMs) have excelled in image generation, offering enhanced stability and quality over methods like GANs [12, 26, 27] and VAEs [30, 57, 42]. Foundational studies [48, 19, 49, 50, 32, 25] established the groundwork for DMs’ capabilities in efficiently scaling up with diverse data. Recent advancements [43, 41, 37, 45, 67, 69, 35] have further improved controllability and user interaction, enabling the creation of images that better reflect user intentions.
Recently, Video Diffusion Models (VDMs) [20] have been proposed for utilizing DMs for video generation tasks. VDMs are capable of generating videos with a wide variety of motions in text-to-video synthesis tasks, supported by the integration of text encoders [40], as shown in [16, 68, 22, 18, 1, 2]. Many open-source text-to-video models have been introduced, including ModelScope[60], AnimateDiff[16], VideoCrafter[6] and so on. These models typically require a pre-trained image generation model, e.g., Stable Diffusion(SD) [44], and introduce additional temporal or motion modules. However, unlike images that contain rich semantic information, text conditions are more difficult to ensure consistency between different frames of the generated video. At the same time, some work also uses image conditions to achieve image-to-video generation with improved spatial semantic consistency [15, 23, 1]. Some works have proposed the paradigm of text-to-image-to-video[11], but image conditions alone cannot effectively control the motion of videos. Combining text and image conditions leads to enhanced spatiotemporal consistency in a text-and-image-to-video workflow [70, 6, 7, 64, 14], but these methods require additional training.
@@
To this end, our research goal in this work is to develop an effective plug-and-play method that is training-free, and can be applied to various text-to-video models to improve the performance of generated videos. To solve this problem, we first attempt to ensure that the semantic information between each frame of the video is consistent in principle. This principle draws inspiration from some previous research in attention-based control [17, 56, 4, 66]. These works have demonstrated in DMs that the queries in attention layers determine the spatial information of the generated image, and correspondingly, values determine the semantic information. We observe that this finding also holds in VDMs and propose the cross-frame self-attention control method. We thus apply the keys and values of the first frame in self-attention layers to each frame and achieve satisfying consistency in the generated video.
Secondly, we observe that as the video’s consistency improves, the motions within videos tend to become less pronounced. To solve this problem, we propose the motion injection mechanism. Based on the assumption that queries control spatial information[4], we divide the sampling process into two branches, namely, an output branch for cross-frame self-attention control, and a motion branch without any attention control. We reserve the queries in the motion branch as motion queries, and use the corresponding motion queries in the output branch. Through the cross-frame self-attention control and motion injection, we ensure that the semantic information between each frame of the video is consistent, while the motion is preserved.
Lastly, we note that motion queries cannot guarantee the spatio-temporal consistency of video. Observing that the output of the output branch has a better spatiotemporal consistency, we further propose the spatiotemporal synchronization, that is, before each sampling step, the latent of the output branch is copied as the initial value of the latent of the motion branch. Our UniCtrl framework combines the above three methods into a plug-in-and-play method that can improve the quality of spatiotemporal consistency and motion quality of the generated videos, while ensuring the consistency of the semantic information of each frame of the video. Through experiments, several text-to-video models have been improved after applying the UniCtrl method, proving the effectiveness and universality of UniCtrl. As illustrated in Fig. 1, UniCtrl plays a significant role in improving spatiotemporal consistency and preserving motion dynamics of generated frames. This method can be readily applied during inference without the need for parameter tuning.
@@ 2 Related Work
Many previous efforts have explored the task of video generation, e.g., GAN-based models [47, 55, 3] and transformer-based models[21, 59, 61, 62]. Recently, following that Diffusion models (DMs) [48, 19, 49, 50, 32, 25] have achieved remarkable results in image generation [43, 41, 37, 45, 36], video diffusion models (VDMs) [20] has also demonstrated their capabilities in video generation [2, 11, 14, 18, 22, 46, 68, 16, 6, 7, 64, 60]. At present, VDMs are mainly implemented by adding additional temporary layers to 2D UNet, which leads to a lack of cross-frame constraints in the training process of the 2D UNet model. Some methods [63, 39, 13] tried to use a training-free method to make the generated videos more smooth. However, how to maintain the cross-frame consistency in videos generated by VDMs remains unresolved.
@@
Different from models that require training[69, 67, 35], attention control [17, 56, 4, 66, 9] is a training-free method which has been widely applied in the task of image editing. Previous work has found that Attention Control can be used to ensure both semantic [4] and spatial consistency [17, 56, 9] in image editing. InfEdit[66] unified the control of semantic consistency and spatial consistency for the first time, proposing unified attention control (UAC). Text2Video-Zero [28] applies frame-level self-attention on text-to-image synthesis methods and enables text-to-video through manipulating motion dynamics in latent codes. Some work has introduced attention control to VDMs for video editing[31, 10, 29], but no one has improved the consistency of generated videos through video diffusion by attention control. Is that possible to introduce UAC into VDMs to ensure cross-frame semantic consistency and spatial consistency throughout the video is an interesting and worthwhile question to explore.
@@ 3.1 Diffusion Models (DMs)
Diffusion Models (DMs) [49, 19, 50, 25] are a type of generative model trained via score matching [24, 33, 51, 52]. The forward process gradually adds noise to data to make it follow the Gaussian distribution: $z_{t}=\mathcal{N}(z_{t};\sqrt{\bar{a}_{t}}z_{0},(1-\sqrt{\bar{a}_{t}})I)$, where $z_{0}$ are samples from the data distribution and $\alpha_{1},\dots,\alpha_{T}$ are from a variance schedule. In [19], this process is re-parameterized into the following form:
@@
The sampling process of DMs[19, 49] is iterative and can be represented in the form with different noise schedule $\sigma_{t}$:
@@
Latent Diffusion Models (LDMs) [43] encodes samples into latents using an encoder $\mathcal{E}$, such that $z_{0}=\mathcal{E}(x_{0})$. Also, the output is reconstructed via a decoder $\mathcal{D}$, represented by $\mathcal{D}(z)$. This approach has led to improvements in stability and efficiency in the training and generation process.
Video Diffusion Models(VDMs)[20] extended the application of DMs to the domain of video generation, adapting the framework to handle 4D video tensors in the form of $frames\times height\times width\times channels$, which we can use $z^{f}_{t}$ to describe the frame $f+1$ at the timestep $t$.
@@ 3.2 Attention Control
We follow the notation from [17, 66]. In the fundamental unit of the diffusion UNet model, there are two main components: cross-attention and self-attention blocks. The process begins with the linear projection of spatial features to generate queries ($Q$). In the self-attention block, both keys ($K$) and values ($V$) are derived from the spatial features through linear projection. Conversely, for the cross-attention part, text features undergo a linear transformation to form keys ($K$) and values ($V$). The attention mechanism [58] can be described as:
$\textsc{Attention}(Q,K,V)={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}M}V={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\textrm{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)}V$ (4)
Mutual Self-Attenion Control (MasaCtrl) is proposed by [4], and find that replacing the $Q$s in attention layers while keeping the $K$s and $V$s same, can change the spatial information of generated images but keeping the semantic information preserved. This technique can help diffusion in spatial-level editing, e.g. a sitting dog to a running dog. Here we use $(\cdot)^{\textrm{src}}$ to represent the tensor obtained from the source image and $(\cdot)^{\textrm{tgt}}$ for the target output we want, we can use the following formula to define the MasaCtrl algorithm.
$\small out=\textsc{Attention}(Q^{\textrm{tgt}},K^{\textrm{src}},V^{\textrm{src}})$
Cross-Attenion Control (P2P) is a method mentioned in [17], which is for semantic-level image editing (e.g. dog to cat). This work observed that different $V$s from different text prompts decide the semantic information of generated images. If attention maps ($M$) in cross-attention layers have been reserved, but use different Vs for the attention calculation, most of the spatial information will be preserved. We here use $(\cdot)^{\textrm{src}}$ to represent the tensor obtained from the source image and prompt and $(\cdot)^{\textrm{tgt}}$ for the target output with target prompt, this algorithm can be described as following as same $Q^{\textrm{src}},K^{\textrm{src}}$ lead to same $M^{\textrm{src}}$:
@@ 4 UniCtrl: Cross-Frame Unified Attention Control
Figure 2: The first row demonstrates the original frames generated with baseline model, which the vehicle and the road are clearly inconsistent across the frames. The second row shows frames generated with baseline model augmented with Cross-Frame Self-Attention Control (SAC). While it maintains incredible spatiotemporary consistency, it exhibits little motion. The third row explains frames augmented with SAC and Motion Injection (MI). Although MI injects more motion in addition to SAC, the results clear demonstates that it falls short on spatiotemporal consistency again. The fourth row contains frames further augmented with Spatiotemporal Synchronization (SS) in addition to SAC and MI, which improves spatiotemporal consistency over the results from the thrid row and achieved balance between motion and spatiotemporary consistency, both in-frame and cross-frame.
In the text-to-video task, it is difficult to ensure the consistency between different frames of the generated video due to the lack of semantic level conditions and the constraint of different frames in the 2D UNet layers. Based on the previous work of DMs[17, 56, 66, 4], it is found that the queries in the attention layers determine the spatial information of the generated images, and correspondingly, the values determine the semantic information. We assume these properties still exist in VDMs.
We first analyzed the role of keys and values in the self-attention layers of VDMs, and then analyzed the role of queries in all the attention layers. Inspired by InfEdit [66], we then propose the cross-frame unified attention control to achieve both semantic level consistency and better spatiotemporal consistency. Lastly, we apply spatiotemporal synchronization by replacing the latent of the motion branch with the latent from the output branch at each sampling step. We demonstate the effectiveness of each module of UniCtrl in Fig. 2.
@@ 4.1 Cross-Frame Self-Attention Control
Previous work [4, 17, 56, 66] has observed that queries in the attention mechanism form the layout and semantic information of generated images[4, 66, 56], while values contribute to the semantic information[17, 66]. Therefore, we hypothesize that using the same values in the attention of different frames can ensure cross-frame consistency. Additionally, we observed that the mismatch between keys and values will degrade the quality of generated videos through our experiment and we provide one qualitative example in Section 5.4. Consequently, in our cross-frame Self-Attention Control (SAC), we inject both keys and values from the first frame, as detailed in Algorithm 1, to every other frame at each self-attention layer during denoising. We showcase one example of the effectiveness of SAC by comparing the first row and the second row in Fig. 2.
@@
3:$z^{0}=z[:,:,0,:,:].\textrm{unsqueeze}(2)$ Extract the first frame.
@@
6:$K^{0}=\textrm{to\_k}(z^{0})$ Obtain the Key of the first frame.
7:$V^{0}=\textrm{to\_v}(z^{0})$ Obtain the Value of the first frame,
@@
4.2 Motion Injection
Figure 3: In our framework, we use key and value from the first frame as represented by $K^{0}$ and $V^{0}$ in the self-attention block. We also use another branch to keep the motion query $Q_{\textrm{m}}$ for motion control. At the beginning of every sampling step, we let the motion latent equal to the sampling latent, to avoid spatial-temporal inconsistency. We explain details of our framework in Algorithm 1 and Algorithm 2.
In our experiments, detailed in Section 5, we identify a significant limitation associated with cross-frame self-attention control: it tends to produce overly similar consecutive frames, leading to minimal motion within the video sequence. Again, as described in [4, 66], the queries in self-attention and cross-attention determine the video’s spatial information, which corresponds to motion. Therefore, by preserving the original queries, we can maintain the motion effect.
As in Fig. 3, we have divided the inference process into two branches: the output branch, which undergoes cross-frame self-attention control, and the motion branch, which does not involve attention control. We retain the queries in the motion branch as motion queries and use the corresponding motion queries in the output branch. Here, we denote motion queries as $Q_{\textrm{m}}$. The method for motion injection can be expressed by the following formula: $out=\textsc{Attention}(Q_{\textrm{m}},\cdot,\cdot)$, where Attention refers to both of self- and cross-attention layers. This method is designed to fully retain the motion of the original video. However, in practice, a trade-off between motion preservation and spatiotemporal coherence remains evident. Consequently, we propose an additional refinement: selectively preserving motion at specific steps throughout the sampling process to enhance control. To this end, we introduce a technique that modulates motion through the integration of a motion injection degree, $c$, which is defined within the interval $0\leq c\leq 1$. The following outlines this approach in detail:
@@ 4.2 Motion Injection
4.3 Spatiotemporal Synchronization
Algorithm 2 Cross-Frame Unified Attention Control
@@ 4.3 Spatiotemporal Synchronization
5:     Controlled Self Attention Algorithm SelfAttnCtrl
6:     Timestep condition for Motion Control $t$
7:$z_{\textrm{m},\tau_{0}}=z_{\tau_{0}}\sim\mathcal{N}(\bm{0},\bm{I})$
[…]
9:for $n=0$ to $N-1$ do
10:     $z_{\textrm{m},\tau_{n}}=z_{\tau_{n}}$
11:     $Q_{\textrm{m}}\leftarrow\textbf{VDM}(z_{\textrm{m},\tau_{n}},c,\tau_{n})$
12:     if $t\geq\tau_{n}$ then
13:         $z_{\tau_{n+1}}=\textbf{VDM}(z_{\tau_{n}},c,\tau_{n})\{Q\leftarrow Q_{\textrm{m}}\}$
14:     else
15:         $z_{\tau_{n+1}}=\textbf{VDM}(z_{\tau_{n}},c,\tau_{n})$
16:     end if
17:end for
18:Output: $z_{\tau_{N}}$

### v4

@@ 1 Introduction
Diffusion Models (DMs) have excelled in image generation, offering enhanced stability and quality over methods like GANs (Goodfellow et al., 2020; Karras et al., 2019; 2020) and VAEs (Kingma & Welling, 2014; Van Den Oord et al., 2017; Ramesh et al., 2021). The superior image generation capability of diffusion models stems from the critical role of the attention mechanism(Rombach et al., 2022a; Peebles & Xie, 2023; Dhariwal & Nichol, 2021). Foundational studies (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020a; 2023; Luo et al., 2023; Karras et al., 2022) established the groundwork for DMs’ capabilities in efficiently scaling up with diverse data. Recent advancements (Rombach et al., 2022a; Ramesh et al., 2022; Nichol et al., 2022; Saharia et al., 2022; Xu et al., 2023b; Zhang et al., 2023b; Mou et al., 2023) have further improved controllability and user interaction, enabling the creation of images that better reflect user intentions.
Recently, Video Diffusion Models (VDMs) (Ho et al., 2022b) have been proposed for utilizing DMs for video generation tasks. VDMs are capable of generating videos with a wide variety of motions in text-to-video synthesis tasks, supported by the integration of text encoders (Radford et al., 2021), self attention, cross attention and temporal attention, as shown in (Guo et al., 2023b; Zhang et al., 2023a; Hu et al., 2022; Ho et al., 2022a; Blattmann et al., 2023a; b). Many open-source text-to-video models have been introduced, including ModelScope(Wang et al., 2023), AnimateDiff(Guo et al., 2023b), VideoCrafter(Chen et al., 2023) and so on. These models typically require a pre-trained image generation model, e.g., Stable Diffusion(SD) (Rombach et al., 2022b), and introduce additional temporal or motion modules. However, unlike images that contain rich semantic information, text conditions are more difficult to ensure consistency between different frames of the generated video. At the same time, some work also uses image conditions to achieve image-to-video generation with improved spatial semantic consistency (Guo et al., 2023a; Hu & Xu, 2023; Blattmann et al., 2023a). Some works have proposed the paradigm of text-to-image-to-video(Girdhar et al., 2023), but image conditions alone cannot effectively control the motion of videos. Combining text and image conditions leads to enhanced spatiotemporal consistency in a text-and-image-to-video workflow (Zhang et al., 2023c; Chen et al., 2023; 2024; Xing et al., 2023; Gu et al., 2023b), but these methods require additional training.
@@
To this end, our research goal in this work is to develop an effective plug-and-play method that is training-free, and can be applied to various text-to-video models to improve the performance of generated videos. To solve this problem, we first attempt to ensure that the semantic information between each frame of the video is consistent in principle. As the attention mechanism plays a significant role, this principle draws inspiration from previous research in attention-based control (Hertz et al., 2023; Tumanyan et al., 2023; Cao et al., 2023; Xu et al., 2023a). These works have demonstrated in DMs that the queries in attention layers determine the spatial information of the generated image, and correspondingly, values determine the semantic information. We observe that this finding also holds in VDMs and propose the cross-frame self-attention control method. We thus apply the keys and values of the first frame in self-attention layers to each frame and achieve satisfying consistency in the generated video.
Secondly, we observe that as the video’s consistency improves, the motions within videos tend to become less pronounced. To solve this problem, we propose the motion injection mechanism. Based on the assumption that queries control spatial information (Cao et al., 2023), we divide the sampling process into two branches: an output branch for cross-frame self-attention control and a motion branch without any attention control. We reserve the queries in the motion branch as motion queries, and use the corresponding motion queries in the output branch. Through the cross-frame self-attention control and motion injection, we ensure that the semantic information between each video frame is consistent, while the motion is preserved.
Lastly, we note that motion queries cannot guarantee the spatiotemporal consistency of video. Observing that the output of the output branch has a better spatiotemporal consistency, we further propose spatiotemporal synchronization, that is, before each sampling step, the latent of the output branch is copied as the initial value of the latent of the motion branch. Our UniCtrl framework combines the above three methods into a plug-in-and-play method that can improve the quality of spatiotemporal consistency and motion quality of the generated videos, while ensuring the consistency of the semantic information of each frame of the video. A recent survey (Melnik et al., 2024) also emphasizes the importance of enhancing spatiotemporal coherence and motion consistency to advance video diffusion models. In the other domains related to video generation, various methods attempt to address similar challenges (Huang et al., 2023; Yang et al., 2023). Through experiments, several text-to-video models have been improved after applying the UniCtrl method, proving the effectiveness and universality of UniCtrl. As illustrated in Figure 1, UniCtrl plays a significant role in improving spatiotemporal consistency and preserving the motion dynamics of generated frames. This method can be readily applied during inference without the need for parameter tuning.
@@ 2 Related Work
Many previous efforts have explored the task of video generation, e.g., GAN-based models (Skorokhodov et al., 2022; Tian et al., 2021; Brooks et al., 2022) and transformer-based models(Hong et al., 2022; Villegas et al., 2022; Wu et al., 2021; 2022). Recently, following that Diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020a; 2023; Luo et al., 2023; Karras et al., 2022) have achieved remarkable results in image generation (Rombach et al., 2022a; Ramesh et al., 2022; Nichol et al., 2022; Saharia et al., 2022; Nichol & Dhariwal, 2021), video diffusion models (VDMs) (Ho et al., 2022b) has also demonstrated their capabilities in video generation (Blattmann et al., 2023b; Girdhar et al., 2023; Gu et al., 2023b; Ho et al., 2022a; Hu et al., 2022; Singer et al., 2022; Zhang et al., 2023a; Guo et al., 2023b; Chen et al., 2023; 2024; Xing et al., 2023; Wang et al., 2023). At present, VDMs are mainly implemented by adding additional temporary layers to 2D UNet, which leads to a lack of cross-frame constraints in the training process of the 2D UNet model. Some methods (Wu et al., 2023; Qiu et al., 2023; Gu et al., 2023a) tried to use a training-free method to make the generated videos more smooth. However, how to maintain the cross-frame consistency in videos generated by VDMs remains unresolved.
@@
Different from models that require training(Zhang et al., 2023b; Xu et al., 2023b; Mou et al., 2023), attention control (Hertz et al., 2023; Tumanyan et al., 2023; Cao et al., 2023; Xu et al., 2023a; Ge et al., 2023b) is a training-free method which has been widely applied in the task of image editing. Previous work has found that Attention Control can be used to ensure both semantic (Cao et al., 2023) and spatial consistency (Hertz et al., 2023; Tumanyan et al., 2023; Ge et al., 2023b) in image editing. InfEdit(Xu et al., 2023a) unified the control of semantic consistency and spatial consistency for the first time, proposing unified attention control (UAC). Text2Video-Zero (Khachatryan et al., 2023) applies frame-level self-attention on text-to-image synthesis methods and enables text-to-video through manipulating motion dynamics in latent codes. Some work has introduced attention control to VDMs for video editing (Liu et al., 2023; Geyer et al., 2023; Khandelwal, 2023), but no one has improved the consistency of generated videos through video diffusion by attention control. Is that possible to introduce UAC into VDMs to ensure cross-frame semantic consistency and spatial consistency throughout the video is an interesting and worthwhile question to explore.
@@ 3.1 Diffusion Models (DMs)
Diffusion Models (DMs) (Song et al., 2020a; Ho et al., 2020; Song et al., 2023; Karras et al., 2022) are a type of generative model trained via score matching (Hyvärinen & Dayan, 2005; Lyu, 2012; Song & Ermon, 2019; Song et al., 2020b). The forward process gradually adds noise to data to make it follow the Gaussian distribution: $z_{t}=\mathcal{N}(z_{t};\sqrt{\bar{a}_{t}}z_{0},(1-\sqrt{\bar{a}_{t}})I)$, where $z_{0}$ are samples from the data distribution and $\alpha_{1},\dots,\alpha_{T}$ are from a variance schedule. In Ho et al. (2020), this process is re-parameterized into the following form:
@@
The sampling process of DMs (Ho et al., 2020; Song et al., 2020a) is iterative and can be represented in the form with different noise schedule $\sigma_{t}$:
@@
Latent Diffusion Models (LDMs) (Rombach et al., 2022a) encodes samples into latents using an encoder $\mathcal{E}$, such that $z_{0}=\mathcal{E}(x_{0})$. Also, the output is reconstructed via a decoder $\mathcal{D}$, represented by $\mathcal{D}(z)$. This approach has led to improvements in stability and efficiency in the training and generation process.
Video Diffusion Models(VDMs) (Ho et al., 2022b) extended the application of DMs to the domain of video generation, adapting the framework to handle 4D video tensors in the form of $frames\times height\times width\times channels$, which we can use $z^{f}_{t}$ to describe the frame $f+1$ at the timestep $t$.
@@ 3.2 Attention Control
We follow the notation from (Hertz et al., 2023; Xu et al., 2023a). In the fundamental unit of the diffusion UNet model, there are two main components: cross-attention and self-attention blocks. The process begins with the linear projection of spatial features to generate queries ($Q$). In the self-attention block, both keys ($K$) and values ($V$) are derived from the spatial features through linear projection. Conversely, for the cross-attention part, text features undergo a linear transformation to form keys ($K$) and values ($V$). The attention mechanism (Vaswani et al., 2017) can be described as:
$\textsc{Attention}(Q,K,V)={\color[rgb]{1,0,0}M}V={\color[rgb]{1,0,0}\textrm{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)}V$ (4)
Mutual Self-Attenion Control (MasaCtrl) is proposed by (Cao et al., 2023), and find that replacing the $Q$s in attention layers while keeping the $K$s and $V$s same, can change the spatial information of generated images but keeping the semantic information preserved. This technique can help diffusion in spatial-level editing, e.g. a sitting dog to a running dog. Here we use $(\cdot)^{\textrm{src}}$ to represent the tensor obtained from the source image and $(\cdot)^{\textrm{tgt}}$ for the target output we want, we can use the following formula to define the MasaCtrl algorithm.
$out=\textsc{Attention}(Q^{\textrm{tgt}},K^{\textrm{src}},V^{\textrm{src}})$
Cross-Attenion Control (P2P) is a method mentioned in (Hertz et al., 2023), which is for semantic-level image editing (e.g. dog to cat). This work observed that different $V$s from different text prompts decide the semantic information of generated images. If attention maps ($M$) in cross-attention layers have been reserved, but use different Vs for the attention calculation, most of the spatial information will be preserved. We here use $(\cdot)^{\textrm{src}}$ to represent the tensor obtained from the source image and prompt and $(\cdot)^{\textrm{tgt}}$ for the target output with target prompt, this algorithm can be described as following as same $Q^{\textrm{src}},K^{\textrm{src}}$ lead to same $M^{\textrm{src}}$:
@@ 4 UniCtrl: Cross-Frame Unified Attention Control
Figure 2: The first row demonstrates the original frames generated with the baseline model, in which the vehicle and the road are inconsistent across the frames. The second row shows frames generated with baseline model augmented with cross-frame Self-Attention Control (SAC). While it maintains incredible spatiotemporal consistency, it exhibits little motion. The third row explains frames augmented with SAC and Motion Injection (MI). Although MI injects more motion in addition to SAC, the results demonstrate that it falls short on spatiotemporal consistency again. The fourth row contains frames further augmented with Spatiotemporal Synchronization (SS) in addition to SAC and MI, which improves spatiotemporal consistency over the results from the third row and achieves a balance between motion and spatiotemporal consistency, both in-frame and cross-frame.
In the text-to-video task, it is difficult to ensure the consistency between different frames of the generated video due to the lack of semantic level conditions and the constraint of different frames in the 2D UNet layers. Based on the previous work of DMs(Hertz et al., 2023; Tumanyan et al., 2023; Xu et al., 2023a; Cao et al., 2023), it is found that the queries in the attention layers determine the spatial information of the generated images, and correspondingly, the values determine the semantic information. We assume these properties still exist in VDMs.
We first analyzed the role of keys and values in the self-attention layers of VDMs, and then analyzed the role of queries in all the attention layers. Inspired by InfEdit (Xu et al., 2023a), we then propose the cross-frame unified attention control to achieve both semantic level consistency and better spatiotemporal consistency. Lastly, we apply Spatiotemporal Synchronization (SS) by replacing the latent of the motion branch with the latent from the output branch at each sampling step. We demonstate the effectiveness of each module of UniCtrl in Figure 2.
@@ 4.1 Cross-Frame Self-Attention Control
Previous work (Cao et al., 2023; Hertz et al., 2023; Tumanyan et al., 2023; Xu et al., 2023a) has observed that queries in the attention mechanism form the layout and semantic information of generated images(Cao et al., 2023; Xu et al., 2023a; Tumanyan et al., 2023), while values contribute to the semantic information(Hertz et al., 2023; Xu et al., 2023a). Therefore, we hypothesize that using the same values in the attention of different frames can ensure cross-frame consistency. Additionally, we observed that the mismatch between keys and values will degrade the quality of generated videos through our experiment and we provide one qualitative example in Section 5.4. Consequently, in our cross-frame Self-Attention Control (SAC), we inject both keys and values from the first frame, as detailed in Algorithm 1, to every other frame at each self-attention layer during denoising. We showcase one example of the effectiveness of SAC by comparing the first row and the second row in Figure 2.
4.2 Motion Injection
Figure 3: In our framework, we use key and value from the first frame as represented by $K^{0}$ and $V^{0}$ in the self-attention block. We also use another branch to keep the motion query $Q_{\textrm{m}}$ for motion control. At the beginning of every sampling step, we let the motion latent equal to the sampling latent, to avoid spatial-temporal inconsistency. Note that in the actual workflow, Q replacement occurs only in cross-attention, as the Q in the self-attention blocks of both branches are always the same. We explain details of our framework in Algorithm 1 and Algorithm 2.
@@ 4.2 Motion Injection
3:$z^{0}=z[:,:,0,:,:].\textrm{unsqueeze}(2)$ Get the 1st frame.
@@
6:$K^{0}=\textrm{to\_k}(z^{0})$ Get the Key of the first frame.
7:$V^{0}=\textrm{to\_v}(z^{0})$ Get the Value of the first frame.
@@
In our experiments, detailed in Section 5, we identify a significant limitation associated with cross-frame self-attention control: it tends to produce overly similar consecutive frames, leading to minimal motion within the video sequence. Again, as described in (Cao et al., 2023; Xu et al., 2023a), the queries in self-attention and cross-attention determine the video’s spatial information, which corresponds to motion. Therefore, by preserving the original queries, we can maintain the motion effect.
As in Figure 3, we have divided the inference process into two branches: the output branch, which undergoes cross-frame self-attention control, and the motion branch, which does not involve attention control. We retain the queries in the motion branch as motion queries and use the corresponding motion queries in the output branch. Here, we denote motion queries as $Q_{\textrm{m}}$. The method for motion injection can be expressed by the following formula: $out=\textsc{Attention}(Q_{\textrm{m}},\cdot,\cdot)$, where Attention refers to both of self- and cross-attention layers. This method is designed to fully retain the motion of the original video. However, in practice, a trade-off between motion preservation and spatiotemporal coherence remains evident. Consequently, we propose an additional refinement: selectively preserving motion at specific steps throughout the sampling process to enhance control. To this end, we introduce a technique that modulates motion through the integration of a motion injection degree, $c$, which is defined within the interval $0\leq c\leq 1$. This approach is further detailed and formalized presented in Algorithm 2 and the following outlines this approach in detail :
@@
Algorithm 2 Motion Injection
@@
5:     Timestep condition for Motion Control $t$
6:$z_{\textrm{m},\tau_{0}}=z_{\tau_{0}}\sim\mathcal{N}(\bm{0},\bm{I})$
7:for $n=0$ to $N-1$ do
8:     $z_{\textrm{m},\tau_{n}}=z_{\tau_{n}}$
9:     $Q_{\textrm{m}}\leftarrow\textbf{VDM}(z_{\textrm{m},\tau_{n}},c,\tau_{n})$
10:     if $t\geq\tau_{n}$ then
11:         $z_{\tau_{n+1}}=\textbf{VDM}(z_{\tau_{n}},c,\tau_{n})\{Q\leftarrow Q_{\textrm{m}}\}$
12:     else
13:         $z_{\tau_{n+1}}=\textbf{VDM}(z_{\tau_{n}},c,\tau_{n})$
14:     end if
15:end for
16:Output: $z_{\tau_{N}}$
4.3 Spatiotemporal Synchronization